For a particular reaction at 128.7 °C, deltaG = – 784.97 kJ/mol, and deltaS = 676.82 J/(mol · K). Calculate deltaG for this reaction at - 77.1 °C.
Answer #1

As we know DG = DH- TIS At T = 128.7°C C OG = - 784.97 KJ/mol D OG =- 784.97 kJ/mol AS=676.82 J mobk). = 401-85 K = 0.67682 k